# Ordered random walks and the Airy line ensemble

### v1

@@ 5.3 Convergence to the Airy line ensemble
Let $r^{T}_{L}=T+2T^{1-a/3}L$, let $\rho_{T}:=\inf\{t\geq r^{T}_{L}:S(t)\in W_{T,\varepsilon}\}$ and $\mathcal{A}_{2T}=\{\rho_{T}\leq 2T\}$. Let $\mathcal{C}_{2T}:=\cap_{j=1}^{d}\mathcal{C}_{j,2T}$ where
$\mathcal{C}_{j,2T}:=\left\{\sup_{0\leq t\leq\rho_{T}\wedge 2T}|B_{j}(t)-S_{j}(t)|\leq T^{\beta}\right\}.$
Suppose that $f:D[-L,L]\rightarrow[0,1]$ is uniformly continuous. We also suppose that $f$ depends only on the co-ordinates $(X^{T}_{d-k+1},\ldots,X^{T}_{d})$; this fact is only used in one place, to establish (17), and can be replaced with other conditions on $f$.
@@
By Lemma 16 the second sum is $o(1)$ for $\frac{\beta}{2}>a$. Next the first summand, by harmonicity of $V$ can be written as
$\mathbf{E}_{x}[f(X^{T})V(S(2T));\tau>2T,\mathcal{C}_{2T}]$
$=\mathbf{E}_{x}[f(X^{T})V(S(\rho_{T}\wedge 2T));\tau>\rho_{T}\wedge 2T,\mathcal{C}_{2T}]$
$=\mathbf{E}_{x}[f(X^{T})V(S(\rho_{T}));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}]$
$\hskip 28.45274pt+\mathbf{E}_{x}[f(X^{T})V(S(2T));\tau>2T,\mathcal{A}^{c}_{2T},\mathcal{C}_{2T}].$
By Lemma 17 the second term is again $o(1)$ for $4a<\varepsilon$. As a result
$\mathbf{E}_{x}^{(V)}[f(X^{T})]=\mathbf{E}_{x}[f(X^{T})V(S(\rho_{T}));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}]+o(1).$
The function $f$ is uniformly continuous and depends on the first $k$ co-ordinates so for some $C>0$ on the event $\mathcal{C}_{2T}$,
@@
Therefore using this bound and martingale properties
$\frac{1}{V(x)}\mathbf{E}_{x}\left[(f(X^{T})-f(Y^{T}))V(S(\rho_{T}));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}\right]$
$\leq CT^{\beta-\gamma}\frac{1}{V(x)}\mathbf{E}_{x}\left[V(S(\rho_{T}));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}\right]$
@@
$M_{1}=\frac{1}{V(x)}\mathbf{E}_{x}[f(Y^{T})V(S(\rho_{T})));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}].$
On the event $\mathcal{A}_{2T}$ we have $\rho_{T}\leq 2T$. On the event $\mathcal{C}_{2T}$ the following inequalities hold: for all $0\leq t\leq 2T,$
@@
$x^{\pm}=(x_{1},x_{2}\pm 2T^{\beta},\ldots,x_{d}\pm 2(d-1)T^{\beta}).$
Note $x^{-}\in W_{T,\varepsilon}$ if $\beta<1/2-\varepsilon$ since $x\in W_{T,\varepsilon}$ for sufficiently large $T$. Therefore
$M_{1}$ $=\frac{1}{V(x)}\mathbf{E}_{x}[f(Y^{T})V(S(\rho_{T}));\tau>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}]$
$\leq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x^{+}}[f(Y^{T})V(S(\rho_{T}));\tau^{\text{bm}}>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}].$
By definition $S(\rho_{T})\in W_{T,\varepsilon}$ and $V\sim\Delta$ on $W_{T,\varepsilon}$ for $22a<\varepsilon-1/$ 2 by Proposition 3. Therefore, using also (18),
@@
$M_{1}$ $\leq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x^{+}}[f(Y^{T})\Delta(B(\rho_{T})));\tau^{\text{bm}}>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}].$
The martingale property of $\Delta$ for Brownian motion killed when the co-ordinates become disordered can be applied to give
$M_{1}$ $\leq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x^{+}}[f(Y^{T})\Delta(B(\rho_{T})));\tau^{\text{bm}}>\rho_{T},\mathcal{A}_{2T},\mathcal{C}_{2T}]$
$=(1+o(1))\frac{\Delta(x^{+})}{V(x)}\mathbf{E}_{x^{+}}^{(\Delta)}[f(Y^{T});\mathcal{A}_{2T},\mathcal{C}_{2T}].$
Using $V(x)=\Delta(x)(1+o(1))$ for $22a<\varepsilon-1/$ 2 and $\Delta(x)=\Delta(x^{+})(1+o(1))$ for $2a+\beta<1/2-\varepsilon,$
$M_{1}$ $\leq(1+o(1))\mathbf{E}_{x^{+}}^{(\Delta)}[f(Y^{T});\mathcal{A}_{2T},\mathcal{C}_{2T}]$
$=(1+o(1))\mathbf{E}_{x^{+}}^{(\Delta)}[f(Y^{T})].$
To obtain the second line we have used the second equations in Lemma 16 and 17 for $\frac{\beta}{2}>a$ and $\varepsilon>2a$. Using Lemma 19
$\mathbf{E}_{x^{+}}^{(\Delta)}[f(Y^{T})]=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})]$
for $\lvert x_{1}\rvert+\lvert x_{d}\rvert=o(T^{\gamma}).$ In a similar manner we can show that
$M_{1}\geq(1+o(1))\mathbf{E}_{x^{-}}^{(\Delta)}[f(Y^{T})]=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})].$
@@
Then
$\mathbf{E}_{x}^{(V)}[f(X^{T})]=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}(f(X^{T}));\nu_{T}\leq T^{1-\varepsilon}]+\mathbf{E}_{x}^{(V)}[f(X^{T});\nu_{T}\geq T^{1-\varepsilon}].$
@@
$\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T})];\nu_{T}\leq T^{1-\varepsilon}]$
$=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$
@@
$E_{5}$ $\leq\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{n})}^{(V)}[f(X^{T})];\nu_{T}\leq T^{1-\epsilon},\lvert M(\nu_{T})\rvert>T^{\gamma-\delta}/2]$
$E_{6}$ $\leq\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert I(\nu_{T})\rvert>T^{\gamma-\delta}/2].$
Lemma 18 shows that $E_{5}\rightarrow 0$ and a symmetric argument shows that $\qquad E_{6}\rightarrow 0$ for $2\gamma-1+\varepsilon>4a$ where $\delta>0$ is chosen sufficiently small.
As a result
@@
$=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]+o(1)$
$=(1+o(1))\mathbf{E}_{x}^{(V)}[\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$
$=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})],$
where in the second line we use the approximation of ordered random walks and non-intersecting Brownian motion for starting positions in $W_{T,\varepsilon}$ satisfying the bound $\lvert S_{d}(\nu_{T})\rvert+\lvert S_{1}(\nu_{T})\rvert\leq T^{\gamma-\delta}$. In the final line we use again Lemma 17 and Lemma 18.
The conditions needed are $\frac{\beta}{2}>a$, $\varepsilon>4a$, $\beta<\gamma$, $2a<1/2-\varepsilon$, $2a+\beta<1/2-\varepsilon$ and $2\gamma-1+\varepsilon>4a$. Here $\gamma=1/2-a/6$ leads to these conditions being satisfiable if $a<\frac{3}{50}$. Indeed in this case there exists $\beta>\frac{3}{25}$ such that when $\varepsilon=\frac{13}{50}$ all above inequalities are met. ∎
@@ Appendix A Likelihood ratio ordering
As $(-X|X<0)$ is positive and $(X-\theta|X>\theta)$ also has a log-concave density then taking an expectation over (19),
@@
We can repeat the argument with $X$ replaced by $(-X)$ to obtain the statement. ∎
@@
+$\qquad\qquad\qquad\qquad\qquad+\int\int_{\{x>y\}}\phi(x,y)\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)$
=$=\int\int_{\{y>x\}}\phi(x,y)\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)+\phi(y,x)\mathbf{P}(U\in dy)\mathbf{P}(V\in dx)$
$\geq\int\int_{\{y>x\}}\phi(x,y)\big{(}\mathbf{P}(U\in dy)\mathbf{P}(V\in dx)-\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)\big{)}$

### v2

@@ 5.3 Convergence to the Airy line ensemble
To establish weak convergence, it is sufficient to prove convergence of the expectation of any uniformly continuous measurable functional $f:D(k,[-L,L])\rightarrow[0,1].$
@@
Suppose $\beta>0$. We will introduce further conditions on $\beta$ later in the proof. Let $\mathcal{C}_{2T}:=\cap_{j=1}^{d}\mathcal{C}_{j,2T}$ where
$\mathcal{C}_{j,2T}:=\left\{\sup_{0\leq t\leq 2T}|B_{j}(t)-S_{j}(t)|\leq T^{\beta}\right\}.$
@@
By Lemma 16 the second sum is $o(1)$ for $\frac{\beta}{2}>a$. Let $\mathcal{A}_{2T}=\{S(2T)\in W_{T,\varepsilon}\}$ and note that
$\mathbf{E}_{x}[f(X^{T})V(S(2T));\tau>2T,\mathcal{C}_{2T}]\geq\mathbf{E}_{x}[f(X^{T})V(S(2T));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}].$
The function $f:D(k,[-L,L])\rightarrow[0,1]$ is uniformly continuous so for some $C>0$ on the event $\mathcal{C}_{2T}$,
@@
Therefore, using this bound and martingale properties
$\frac{1}{V(x)}\mathbf{E}_{x}\left[(f(X^{T})-f(Y^{T}))V(S(2T));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}\right]$
$\leq CT^{\beta-\gamma}\frac{1}{V(x)}\mathbf{E}_{x}\left[V(S(2T));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}\right]$
@@
$M_{1}=\frac{1}{V(x)}\mathbf{E}_{x}[f(Y^{T})V(S(2T)));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}].$
On the event $\mathcal{C}_{2T}$ the following inequalities hold: for all $0\leq t\leq 2T,$
@@
$x^{-}=(x_{1},x_{2}-2T^{\beta},\ldots,x_{d}-2(d-1)T^{\beta}).$ (23)
Note, for sufficiently large $T$, then $x^{-}\in W_{T,\varepsilon}$ if $\beta<1/2-\varepsilon$ since $x\in W_{T,\varepsilon}$. We now note that we can construct a coupled process $Y^{T}_{-}$ driven by the same Brownian motions, where $Y^{T}_{-}$ has the same definition as $Y^{T}$ but with the Brownian motions started from the initial condition $x^{-}$ instead of $x$. Under the law $\mathbf{E}_{x}$, note that $Y^{T}_{-}$ involves Brownian motions started from $x^{-}$. The coupled systems $Y^{T}$ and $Y^{T}_{-}$ satisfy an analogue of (21) so that by following the argument immediately after (21) we can establish
$\frac{1}{V(x)}\mathbf{E}_{x}[\lvert f(Y^{T})-f(Y^{T}_{-})\rvert V(S(2T));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}]\rightarrow 0.$
Therefore,
$M_{1}$ $=\frac{(1+o(1))}{V(x)}\mathbf{E}_{x}[f(Y^{T}_{-})V(S(2T));\tau>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}].$
Let $\tau^{\text{bm}}_{-}$ denote the first exit time from the Weyl chamber for the Brownian motion started from $x^{-}$. Due to the coupling, we have that $\{\tau^{\text{bm}}_{-}>2T\}\subset\{\tau>2T\}$. As a result,
$M_{1}\geq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x}[f(Y^{T}_{-})V(S(2T));\tau^{\text{bm}}_{-}>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}].$
By definition, when $\mathcal{A}_{2T}$ holds, we have $S(2T)\in W_{T,\varepsilon}$ and $V\sim\Delta$ on $W_{T,\varepsilon}$ for $2a<1/2-\varepsilon$ by Proposition 3. Furthermore, $\Delta(S(2T))=\Delta(B_{-}(2T))(1+o(1))$ for $2a+\beta<1/2-\varepsilon$ by (22), where the Brownian motions $B_{-}$ are started from $x^{-}$. Therefore,
$M_{1}$ $\geq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x^{-}}[f(Y^{T})\Delta(B(2T)));\tau^{\text{bm}}>2T,\mathcal{A}_{2T},\mathcal{C}_{2T}].$
Let $\mathcal{A}_{2T}^{\text{bm}}=\{\frac{1}{2}B(2T)\in W_{T,\varepsilon}\}$. On the event $\mathcal{C}_{2T}$, we have that $\mathcal{A}_{2T}^{\text{bm}}\subset\mathcal{A}_{2T}$ for sufficiently large $T$, due to the factor $1/2$ appearing in the definition of $\mathcal{A}_{2T}^{\text{bm}}$ and that $\beta<1/2-\varepsilon$. Therefore,
$M_{1}$ $\geq\frac{(1+o(1))}{V(x)}\mathbf{E}_{x^{-}}[f(Y^{T})\Delta(B(2T));\tau^{\text{bm}}>2T,\mathcal{A}_{2T}^{\text{bm}},\mathcal{C}_{2T}]$
$=(1+o(1))\frac{\Delta(x^{-})}{V(x)}\mathbf{E}_{x^{-}}\left[f(Y^{T})\frac{\Delta(B(2T))}{\Delta(x^{-})};\tau^{\text{bm}}>2T,\mathcal{A}_{2T}^{\text{bm}},\mathcal{C}_{2T}\right].$
Using $V(x)=\Delta(x)(1+o(1))$ for $2a<1/2-\varepsilon$ and $\Delta(x)=\Delta(x^{-})(1+o(1))$ for $2a+\beta<1/2-\varepsilon,$
$M_{1}$ $\geq(1+o(1))\mathbf{E}_{x^{-}}\left[f(Y^{T})\frac{\Delta(B(2T))}{\Delta(x^{-})};\tau^{\text{bm}}>2T,\mathcal{A}_{2T}^{\text{bm}},\mathcal{C}_{2T}\right]$
$=(1+o(1))\mathbf{E}_{x^{-}}\left[f(Y^{T})\frac{\Delta(B(2T))}{\Delta(x^{-})};\tau^{\text{bm}}>2T,\mathcal{A}_{2T}^{\text{bm}}\right].$
To obtain the second line, we use part (ii) of Lemma 16 for $\frac{\beta}{2}>a$. We now use Corollary 1.5 of [7], which proves that in the Gaussian Unitary Ensemble of size $d\times d$, the minimal gap between eigenvalues multiplied by $d^{4/3}$ converges to a limiting distribution as $d$ grows to infinity. Hence, with probability tending to one, Dyson Brownian motion started from zero is in $W_{T,\varepsilon}$ at time $2T$ for any $\varepsilon>5a/6.$ This property can then be extended to Dyson Brownian motion started from $x^{-}$ using the monotonicity of gaps in part (ii) of Lemma 20. Hence, for $\varepsilon>5a/6$,
@@
$\mathbf{E}_{x^{-}}\left[f(Y^{T})\frac{\Delta(B(2T))}{\Delta(x^{-})};\tau^{\text{bm}}>2T,(\mathcal{A}_{2T}^{\text{bm}})^{c}\right]\rightarrow 0,\quad T\rightarrow\infty.$
Therefore, now using Lemma 19,
$M_{1}$ $\geq(1+o(1))\mathbf{E}_{x^{-}}\left[f(Y^{T})\frac{\Delta(B(2T))}{\Delta(x^{-})};\tau^{\text{bm}}>2T\right]$
$=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})]$
for $x\in W_{T,\varepsilon}$ and $\lvert x_{1}\rvert+\lvert x_{d}\rvert=o(T^{\gamma}).$ Therefore
$\mathbf{E}^{(V)}_{x}[f(X^{T})]\geq(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})].$
Applying the same argument to the function 1f1-f gives
$\mathbf{E}^{(V)}_{x}[(1-f)(X^{T})]\geq(1+o(1))\mathbf{E}_{0}^{(\Delta)}[(1-f)(Y^{T})]$
from which we deduce the equality
$\mathbf{E}^{(V)}_{x}[f(X^{T})]=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})].$
@@
Then by the strong Markov property,
$\mathbf{E}_{x}^{(V)}[f(X^{T})]$ $=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon}]$
$\quad+\mathbf{E}_{x}^{(V)}[f(X^{T});\nu_{T}\geq T^{1-\varepsilon}].$
@@
$\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon}]$
$=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$
@@
$E_{5}$ $\leq\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\epsilon},\lvert M(\nu_{T})\rvert>T^{\gamma-\delta}/2]$
$E_{6}$ $\leq\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon},\lvert I(\nu_{T})\rvert>T^{\gamma-\delta}/2].$
Lemma 18 shows that $E_{5}\rightarrow 0$ and $E_{6}\rightarrow 0$ for $2\gamma-1+\varepsilon>4a$ where $\delta>0$ is chosen sufficiently small.
As a result,
@@
$=\mathbf{E}_{x}^{(V)}[\mathbf{E}_{S(\nu_{T})}^{(V)}[f(X^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]+o(1)$
$=(1+o(1))\mathbf{E}_{x}^{(V)}[\mathbf{E}_{0}^{(\Delta)}[f(Y^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$
where in the second line we use the approximation of ordered random walks and non-intersecting Brownian motion for starting positions in $W_{T,\varepsilon}$ satisfying the bound $\lvert S_{d}(\nu_{T})\rvert+\lvert S_{1}(\nu_{T})\rvert\leq T^{\gamma-\delta}$. We now show that
$\mathbf{E}_{x}^{(V)}[\mathbf{E}_{0}^{(\Delta)}[f(Y^{T}(\cdot-\nu_{T}))];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$ (24)
$=(1+o(1))\mathbf{E}_{x}^{(V)}[\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}].$
For $\gamma>1/2-\varepsilon/2$ and on the event $\{\nu_{T}\leq T^{1-\varepsilon}\}$ we can use uniform continuity of $f$ to obtain that there exists $0<\delta^{\prime}<\gamma-1/2+\varepsilon/2$ such that
$\lvert f(Y^{T}(\cdot-\nu_{T}))-f(Y^{T})\rvert$
$\leq CT^{-\gamma}\sup_{(j,t,s)\in J}\lvert B_{j}(T+2T^{1-a/3}t-s)-B_{j}(T+2T^{1-a/3}t)\rvert$
$\leq CT^{-\gamma+1/2-\varepsilon/2+\delta^{\prime}}\rightarrow 0$
where $J=\{(j,t,s):d-k+1\leq j\leq d,-L\leq t\leq L,0\leq s\leq T^{1-\varepsilon}\}$. This bound is the analogue of (21). Hence, (24) can be established by following the argument immediately after (21). Finally, we use again Lemma 17 and Lemma 18 to obtain,
$\mathbf{E}_{x}^{(V)}[\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})];\nu_{T}\leq T^{1-\varepsilon},\lvert M(\nu_{T})\rvert+\lvert I(\nu_{T})\rvert\leq T^{\gamma-\delta}]$
$=(1+o(1))\mathbf{E}_{0}^{(\Delta)}[f(Y^{T})].$
The conditions needed are $\frac{\beta}{2}>a$, $\varepsilon>4a$, $\beta<\gamma$, $2a<1/2-\varepsilon$, $2a+\beta<1/2-\varepsilon$ and $2\gamma-1+\varepsilon>4a$. Here, $\gamma=1/2-a/6$ leads to these conditions being satisfiable if $a<\frac{3}{50}$. Indeed in this case there exists $\beta>\frac{3}{25}$ such that when $\varepsilon=\frac{13}{50}$ all above inequalities are met. ∎
@@ Appendix A Likelihood ratio ordering
As (X|X<0)(-X|X<0) is positive and $(X-\theta|X>\theta)$ also has a log-concave density then taking an expectation over (25),
@@
We can repeat the argument with $X$ replaced by (X)(-X) to obtain the statement. ∎
@@
$\qquad\qquad\qquad\qquad\qquad+\int\int_{\{x>y\}}\phi(x,y)\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)$
$=\int\int_{\{y>x\}}\phi(x,y)\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)+\phi(y,x)\mathbf{P}(U\in dy)\mathbf{P}(V\in dx)$
$\geq\int\int_{\{y>x\}}\phi(x,y)\big(\mathbf{P}(U\in dx)\mathbf{P}(V\in dy)-\mathbf{P}(U\in dy)\mathbf{P}(V\in dx)\big)$